Find f∘g and g∘f , where f(x)=x^2+2x+1 and g(x)=x^2-20, are functions from R to R.

Type/Insert your answer here!


1
Expert's answer
2021-05-07T09:41:41-0400

"f(x)=x^2+2x+1"

"g(x)=x^2-20"


"f\circ{g}=f(x^2-20)=(x^2-20)^2+2(x^2-20)+1="

"=x^4-40x^2+400+2x^2-40+1="

"=x^4-38x^2+361"


"g\circ{f}=g(x^2+2x+1)="

"=(x^2+2x+1)^2-20="

"=x^4+2x^3+x^2+2x^3+4x^2+2x+x^2+2x+1-20="

"=x^4+4x^3+6x^2+4x-19"